OpenStudy (anonymous):

Solve d - (-2) = 19 for d. d =

OpenStudy (anonymous):

that is the same as d+2 = 19 so subtract 2 from both sides

OpenStudy (nirmalnema):

d=17

OpenStudy (anonymous):

17 or 12

OpenStudy (anonymous):

Solve -5w = -80 for w=

OpenStudy (nirmalnema):

- sign can be cancelled out from both the sides you will get 5w=80 w=80/5 w=16